OpenStudy (anonymous):

HELP plz!!! This table represents the boiling point of water at different altitudes. Which equation represents water’s boiling point (y) based on the change in altitude (x)? Altitude (thousands of feet) Boiling Point of Water (Fahrenheit) (sea level) 0 212 0.5 211.1 1.0 210 2.0 208.4 2.5 207.5 3.0 206.6 4.0 204.8 4.5 203.9 A. -0.6x + 212 B. -1.8x + 212 C. -0.6x + 203.9 D. -1.8x HELP plz

jimthompson5910 (jim_thompson5910):

which altitude would you like to pick?

OpenStudy (anonymous):

any

jimthompson5910 (jim_thompson5910):

let's pick 1.0 or just 1

jimthompson5910 (jim_thompson5910):

so x = 1

jimthompson5910 (jim_thompson5910):

now let's plug this x value into each equation to see which one gives the result of 210

jimthompson5910 (jim_thompson5910):

choice A y = -0.6x + 212 y = -0.6*1 + 212 y = 211.4 So choice A is NOT the answer

jimthompson5910 (jim_thompson5910):

choice B y = -1.8x + 212 y = -1.8*1 + 212 y = 210.2 choice B is NOT the answer

jimthompson5910 (jim_thompson5910):

see how this all works?

OpenStudy (anonymous):

yes sir

jimthompson5910 (jim_thompson5910):

alright that's great so what's the answer?

OpenStudy (anonymous):

i think its C right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you dude

jimthompson5910 (jim_thompson5910):

you're welcome